### v1

# Universal Transient Stability Analysis: A Large Language Model-Enabled Dynamics Prediction Framework

Abstract
Existing dynamics prediction frameworks for transient stability analysis (TSA) fail to achieve multi-scenario ”universality”—the inherent ability of a single, pre-trained architecture to generalize across diverse operating conditions, unseen faults, and heterogeneous systems. To address this, this paper proposes TSA-LLM, a large language model (LLM)-based universal framework that models multi-variate transient dynamics prediction as a univariate generative task with three key innovations: First, a novel data processing pipeline featuring channel independence decomposition to resolve dimensional heterogeneity, sample-wise normalization to eliminate separate stable/unstable pipelines, and temporal patching for efficient long-sequence modeling; Second, a parameter-efficient freeze-and-finetune strategy that augments the LLM’s architecture with dedicated input embedding and output projection layers while freezing core transformer blocks to preserve generic feature extraction capabilities; Third, a two-stage fine-tuning scheme that combines teacher forcing, which feeds the model ground-truth data during initial training, with scheduled sampling, which gradually shifts to leveraging model-generated predictions, to mitigate cumulative errors in long-horizon iterative prediction. Comprehensive testing demonstrates the framework’s universality, as TSA-LLM trained solely on the New England 39-bus system achieves zero-shot generalization to mixed stability conditions and unseen faults, and matches expert performance on the larger Iceland 189-bus system with only 5% fine-tuning data. This multi-scenario versatility validates a universal framework that eliminates scenario-specific retraining and achieves scalability via large-scale parameters and cross-scenario training data.
keywords:
Deep learning, data-driven methods, transient stability analysis, dynamic trajectory prediction, pre-trained large language model, iterative prediction.
journal: Applied Energy
[…]
[label1]organization=Collage of Control Science and Engineering, Zhejiang University,postcode=310027, city=Hangzhou, country=China
[…]
[label2]organization=School of Advanced Manufacturing and Robotics,Peking University,postcode=100871, city=Beijing, country=China
1 Introduction
The increasing penetration of renewable energy sources reduces system inertia and elevates operational uncertainty, posing significant challenges to grid stability [shen2025physics]. Consequently, modern power systems are increasingly susceptible to contingencies, demanding more robust analytical frameworks. A key approach is transient stability analysis (TSA), which evaluates the dynamic behavior of synchronous generators (SGs) following disturbances [ye2024use]. TSA is critical for preventive control and emergency response, enabling grid operators to enhance resilience against cascading failures and ensure reliable operation under evolving operating conditions (OCs). Conventionally, the dynamic behavior of power systems is modeled using nonlinear differential-algebraic equations (DAEs), which are numerically solved via time-domain simulation (TDS) methods. While TDS provides high-fidelity temporal resolution, its computational intensity and scalability challenges limit its online practicality [zhao2022structure]. Alternatively, direct methods assess transient stability by analyzing energy-based criteria without explicitly solving DAEs [qiu2022adaptive]. These methods evaluate system stability by comparing post-fault energy to a predefined critical energy threshold. However, their reliance on simplified system models and inherently conservative critical energy estimates restricts their applicability in real-world scenarios[azman2020unified].
The widespread deployment of phasor measurement units (PMUs) has spurred the development of data-driven TSA. These methods learn the mapping between system measurements and stability-related features, enabling rapid online applications such as real-time stability classification [zhou1994application, gomez2010support, Sun2007An, chen2022interpretable, ren2021interpretable, azman2020unified, zhu2023integrated] and stability margin estimation [su2023online, zhu2019hierarchical, zhu2021networked]. Early research in this domain was dominated by classical machine learning. For instance, artificial neural networks were applied to perform rapid transient stability classification [zhou1994application], while support vector machines demonstrated robust post-fault stability prediction using synchrophasor data [gomez2010support]. Concurrently, decision trees were used to extract critical stability indicators through offline feature ranking, facilitating periodic updates to security rules [Sun2007An]. More recent research has shifted toward deep learning (DL) frameworks to improve both prediction accuracy and model interpretability. To enhance explainability, a time-adaptive attention-based gated recurrent unit (GRU) was developed to visualize feature importance [chen2022interpretable], while other work integrated decision tree path length constraints into the DL loss function [ren2021interpretable]. For performance improvement, hybrid architectures combining convolutional neural networks (CNNs) and long short-term memory (LSTMs) were proposed for stability classification and oscillation prediction [azman2020unified]. Beyond binary classification, significant effort has focused on directly predicting stability margins, often derived from the critical clearing time (CCT). Notable methods include an ensemble DL framework with dynamic error correction for CCT estimation [su2023online] and a two-level CNN regression model that predicts margins from 2-D images of system trajectories [zhu2019hierarchical].
Although existing data-driven models provide rapid and accurate stability classification, this binary output offers limited actionable insight for system operators [zhao2022structure, ye2024use]. Consequently, TSA research has shifted toward predicting detailed dynamic trajectories, which are critical for informing corrective actions such as load shedding or out-of-step (OOS) SG tripping [ye2024use]. One prominent approach involves physics-informed neural networks (PINNs), which incorporate physical constraints as regularization terms to predict system dynamics [misyris2020physics]. However, the applicability of early PINN models was often limited to stable OCs. Subsequent work on physics-following neural networks (PFNNs) has sought to improve generalization across both stable and unstable cases by introducing time-varying algebraic parameters and supervised initialization into the PINN framework [shen2025physics]. Other notable strategies include using LSTMs with separate processing pipeline for stable and unstable samples [ye2024use] and leveraging Fourier neural operators (FNOs) to predict transients in the frequency domain, thereby simplifying the learning task [cui2023frequency].
While the methods in [shen2025physics, ye2024use, cui2023frequency] improve prediction accuracy in mixed stable/unstable OCs, they exhibit poor generalization to unseen faults. To address this, the stochastic variational deep kernel regressor (SVDKR)-based framework in [tan2025bayesian] integrates an OOS detector for classifying SGs as stable or unstable. Dedicated SVDKR predictors are trained per stability class, followed by per-SG error correctors to refine unit-level predictions. Although this improves robustness, the parameter complexity and computational overhead hinder its scalability to large systems with numerous SGs. The deep neural representation (DNR) framework was proposed to mitigate these issues [zhao2022structure], using a unified graph attention and GRU architecture to predict dynamics in a single model. This design enhances performance under unseen faults but suffers from significant cumulative error amplification over long prediction horizons. More importantly, both the SVDKR [tan2025bayesian] and DNR [zhao2022structure] frameworks are designed for a single system. They lack the ability to generalize across heterogeneous power systems with different topologies and SG configurations, necessitating laborious retraining for each new system.
As demonstrated in Table 1, existing frameworks suffer from pronounced performance degradation across diverse OCs, faults and system configurations. This shortcoming motivates the pursuit of universality in TSA models, defined as the intrinsic capability of a single pre-trained architecture to generalize robustly across mixed stable/unstable OCs, unseen fault events, and heterogeneous power systems. Universality enables a paradigm shift in TSA, analogous to the versatility of large language models (LLMs) in natural language processing, where a unified model adeptly addresses diverse dynamic trajectory prediction tasks [achiam2023gpt]. By eliminating the need for scenario-specific retraining, a universal TSA framework provides foundational capabilities, most prominently exemplified by cross-system generalization. This feature enables efficient knowledge transfer from source-domain systems to large-scale target grids via few-shot adaptation, thereby minimizing computational simulation and labor costs associated with custom model development for diverse power networks [tu2024powerpm]. Furthermore, universality inherently supports scalability, with performance systematically improving as model parameters and multi-scenario training datasets expand, which ensures adaptability to evolving grid complexities [liu2024timer].
However, achieving this universality requires overcoming three fundamental limitations: 1) Data processing rigidity, where current methods are constrained to fixed input dimensionalities based on specific state variables, limiting scalability and requiring complete retraining for system changes or monitored SGs modifications [zhao2022structure], while frequently necessitating separate processing pipelines for stable/unstable dynamics that introduce computational overhead and misclassification risks [tan2025bayesian, ye2024use]; 2) Architectural constraints, as prevalent recurrent neural network architectures (LSTMs [ye2024use] and GRUs [zhao2022structure]) suffer from vanishing gradient problems that impede long-range dependency modeling and inherently sequential processing that precludes parallel computation, creating critical bottlenecks for model scalability [liu2024timer]; 3) Training and error propagation, where the temporal asymmetry between short observation windows and extended prediction horizons in online TSA, combined with existing training schemes that lack explicit error suppression mechanisms, causes prediction errors to accumulate at each time step during iterative prediction, severely degrading long-term reliability [shen2025physics, zhao2022structure].
To address these challenges, this study proposes a pre-trained LLM-based universal dynamics prediction framework, TSA-LLM, that formulates multi-variate transient dynamics prediction as a univariate generative sequence prediction task. The main contributions are:
  1. 1.
    To the best of the authors’ knowledge, this is the first work to demonstrate universality in TSA through an LLM-based approach, achieving generalization across diverse OCs, unseen faults, and heterogeneous systems by systematically addressing fundamental limitations in data processing, model architecture, and fine-tuning schemes.
  2. 2.
    A novel data processing pipeline is designed and operates sequentially: channel-independence decouples multivariate dynamic sequences into univariate streams, resolving dimensional heterogeneity across different systems; sample-wise normalization standardizes sequence samples via internal statistics, eliminating separate pipelines for stable/unstable OCs; temporal patching tokenizes sequences into semantic patches, enabling efficient long-sequence modeling while capturing high-level temporal patterns.
  3. 3.
    A Generative Pre-trained Transformer (GPT) architecture is devised that incorporates dedicated input embedding and output projection alongside causal attention mechanisms for temporal dependency modeling. The freeze-and-finetune strategy preserves pre-trained sequence feature extraction capability by freezing core transformer blocks (T-blocks) while selectively fine-tuning TSA-specific layers, achieving balance between knowledge retention and task adaptation.
  4. 4.
    A two-stage fine-tuning scheme is proposed to mitigate error propagation during iterative and generative dynamics prediction. The initial teacher forcing (TeaF) stage ensures stable convergence by exclusively feeding ground-truth sequences as input. Subsequently, the scheduled sampling (SchS) stage enhances robustness by increasingly replacing ground-truth inputs with the model’s own predictions, thereby compelling the model to correct for self-generated errors.
Table 1: Universality Comparison of Methods
Method Universality Scenarios
Single
Model
Stable/Unstable
OCs
Unseen
Faults
Heterogeneous
System
PINN [misyris2020physics] $\times$ $\times$ $\times$ $\checkmark$
LSTM [ye2024use] $\checkmark$ $\times$ $\times$ $\times$
PFNN [shen2025physics] $\checkmark$ $\times$ $\times$ $\checkmark$
FNO [cui2023frequency] $\checkmark$ $\times$ $\times$ $\checkmark$
SVDKR [tan2025bayesian] $\checkmark$ $\checkmark$ $\times$ $\times$
DNR [zhao2022structure] $\checkmark$ $\checkmark$ $\times$ $\checkmark$
TSA-LLM $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2 Preliminary and Problem Formulation
2.1 Power System Model
Considering a power system containing $n_{b}$ buses, $n_{g}$ SGs, and $n_{l}$ loads, the DAEs describing the the system’s transient dynamics, can be formulated as follows [shen2025physics]:
$\begin{cases}&\dot{\mathbf{x}}(t)=\mathbf{f}(\mathbf{x}(t),\mathbf{y}(t),\mathbf{u}(t),\mathcal{F}(t),t),\\ &\mathbf{0}=\mathbf{g}(\mathbf{x}(t),\mathbf{y}(t),\mathbf{u}(t),\mathcal{F}(t),t),\end{cases}$ (1)
where $\mathbf{x}(t)\in\mathbb{R}^{n_{\mathbf{x}}}$ is the vector of dynamic state variables (e.g., rotor angles $\delta_{i}$ and angular speed $\omega_{i}$ of $i$-th SG), and $\mathbf{y}(t)\in\mathbb{R}^{n_{\mathbf{y}}}$ is the vector of algebraic variables (e.g., voltage magnitudes $V_{i}$ of $i$-th bus). The vector $\mathbf{u}(t)\in\mathbb{R}^{n_{\mathbf{u}}}$ represents control inputs, such as governor setpoints. The fault set $\mathcal{F}(t)$ denotes the system components experiencing faults. The nonlinear functions $\mathbf{f}(\cdot)$ and $\mathbf{g}(\cdot)$ describe the system dynamics and algebraic constraints, respectively.
2.2 Transformer Attention Mechanism
The transformer architecture has demonstrated remarkable success in sequence modeling tasks due to its parallelizable computations and ability to capture long-range dependencies [vaswani2017attention]. At the core of this architecture is the self-attention mechanism. As illustrated in the bidirectional attention mechanism presented in Figure 1(a), the process begins with an input sequence $\mathbf{X}=[\mathbf{x}(1),\mathbf{x}(2),\ldots,\mathbf{x}(T)]\in\mathbb{R}^{T\times d}$, where $T$ is the sequence length and $d$ is the feature dimension, the bidirectional attention mechanism first transforms the input into three different representations: queries ($\mathbf{Q}$), keys ($\mathbf{K}$), and values ($\mathbf{V}$):
$\mathbf{Q}=\mathbf{X}\mathbf{W}^{Q},\quad\mathbf{K}=\mathbf{X}\mathbf{W}^{K},\quad\mathbf{V}=\mathbf{X}\mathbf{W}^{V},$ (2)
where $\mathbf{W}^{Q}\in\mathbb{R}^{d\times d_{k}}$, $\mathbf{W}^{K}\in\mathbb{R}^{d\times d_{k}}$, and $\mathbf{W}^{V}\in\mathbb{R}^{d\times d_{v}}$ are learnable parameter matrices. The self-attention matrix is then computed as:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left({\mathbf{Q}\mathbf{K}^{T}}/{\sqrt{d_{k}}}\right)\mathbf{V},$ (3)
where ${1}/{\sqrt{d_{k}}}$ is a scaling factor to prevent vanishing gradients during training. In practice, Multi-Head Attention (MHA) is employed to enrich the model’s expressive capability:
$\text{MHA}(\mathbf{X})=\text{Concat}(\text{head}_{1},\text{head}_{2},\ldots,\text{head}_{h})\mathbf{W}^{O},$ (4)
where $\text{head}_{i}=\text{Attention}(\mathbf{X}\mathbf{W}^{Q}_{i},\mathbf{X}\mathbf{W}^{K}_{i},\mathbf{X}\mathbf{W}^{V}_{i})$. $\mathbf{W}^{Q}_{i}\in\mathbb{R}^{d\times d_{k}/h}$, $\mathbf{W}^{K}_{i}\in\mathbb{R}^{d\times d_{k}/h}$, $\mathbf{W}^{V}_{i}\in\mathbb{R}^{d\times d_{v}/h}$ are parameter matrices for the $i$-th attention head, $\mathbf{W}^{O}\in\mathbb{R}^{d_{v}\times d}$ is the output projection matrix, and $h$ is the number of attention heads. By operating on the entire sequence, this bidirectional attention mechanism is well-suited for tasks like sequence translation and classification [vaswani2017attention].
Figure 1: Comparison of attention mechanisms and iterative prediction in TSA-LLM.
3 Methodology
This section details the proposed TSA-LLM, covering its data processing paradigm, the GPT-based architecture, and the two-stage fine-tuning scheme. Figure 2 illustrates the overall architecture and data workflow.
3.1 Data Processing
3.1.1 Channel Independence
The dynamic state variables of a power system are defined as the multivariate sequences:
$\mathbf{x}(t)=[x_{1}(t),x_{2}(t),\dots,x_{n_{x}}(t)]^{T},$ (5)
where each $x_{i}(t)\in\mathbb{R}$ corresponds to a distinct physical channel (e.g., $\delta_{i}$ or $\omega_{i}$). To overcome the fixed input dimensionality and the significant variance between stable and unstable dynamics, channel independence as illustrated in Figure 2(a) decouples the multivariate sequences $\mathbf{x}(t)$ into a set of univariate series $\{x_{i}(t)\}_{i=1}^{n_{x}}$, which are then processed and predicted independently. This design renders the model architecture agnostic to the number of input variables and enables robust handling of diverse dynamics across heterogeneous systems, thereby enhancing both cross-system scalability.
3.1.2 Segmentation and Sample-wise Normalization
Each univariate sequence $x_{i}\in\mathbb{R}^{T}$ is first partitioned into input-target samples via a sliding window of length $L_{\text{sam}}=L_{\text{seq}}+L_{\text{pred}}$ with a unit stride. Each sample contains an input sequence of length $L_{\text{seq}}$ and its corresponding prediction target of length $L_{\text{pred}}$. The input portion, $x_{ij,\text{input}}\in\mathbb{R}^{L_{\text{seq}}}$, is then subjected to sample-wise normalization to ensure numerical stability, particularly for high-magnitude unstable trajectories:
$x_{ij}=({x_{ij,\text{input}}-\mu_{ij}})/{\sigma_{ij}}\in\mathbb{R}^{L_{\text{seq}}},$ (6)
where the mean $\mu_{ij}$ and standard deviation $\sigma_{ij}$ are computed from $x_{ij,\text{input}}$ rather than the entire training set. This sample-wise normalization, as presented in Figure 2(b) not only manages the extreme values of OOS SGs but also obviates the need for separate normalization statistics for stable/unstable SGs, a common practice in other methods [ye2024use]. Crucially, our approach eliminates the requirement for an online stability classification that may introduce both processing delays and the risk of error [tan2025bayesian], thus guaranteeing a more robust and streamlined framework suitable for real-time application.
3.1.3 Temporal Patch
The final data processing step is temporal patch, designed to enhance both feature representation and modeling efficiency. Each normalized sample $x_{ij}$ is partitioned into $P$ overlapping patches of length $L_{p}$ with a stride $S$, yielding a new, more compact representation ${x}^{P}_{ij}\in\mathbb{R}^{P\times L_{p}}$ as illustrated in Figure 2(c). The number of patches $P$ is given by:
$P=\left\lfloor({L_{\text{seq}}-L_{p}})/{S}\right\rfloor+1,$ (7)
where $\lfloor\cdot\rfloor$ is the floor function. This patch-based tokenization enables the model to capture rich, local dynamic patterns (e.g., transient oscillations) that are more informative than isolated time points. Concurrently, by strategically reducing the input sequence length from $L_{\text{seq}}$ to $P$, it is instrumental in mitigating the quadratic complexity ($\mathcal{O}(N^{2})$) of the attention mechanism [vaswani2017attention], thereby ensuring the computational tractability of long-sequence modeling.
Figure 2: Model structure and pipeline of TSA-LLM.
3.2 Augmented GPT Architecture for TSA
Predicting long-horizon power system dynamics directly from a short observation window is an ill-posed problem, as the limited input resolution often leads to large prediction errors [shen2025physics]. A more principled solution is to adopt an iterative, autoregressive paradigm. In this approach, the model generates the future trajectory one step at a time, where each new prediction is conditioned on the sequence of all prior states (both observed and previously generated) [zhao2022structure]. This process is mathematically equivalent to the fundamental task of generative LLMs. It can be formally expressed by factorizing the conditional probability of the future trajectory, which enforces strict temporal causality:
$p\left(\hat{x}_{ij,\text{target}}\mid{x}_{ij,\text{input}}\right)=\prod_{t=L_{\text{scq }}+1}^{L_{\text{sam}}}p\left(x_{ij}(t)\mid x_{ij}(<t)\right).$ (8)
This inherent similarity is our core motivation for introducing an LLM architecture. We therefore adopt the open-source GPT architecture [radford2019language], a model renowned for its strong performance in generative tasks within natural language processing field. The model’s strength lies in its causal attention mechanism, which is fundamentally designed for the conditional, step-by-step sequence generation that accurate long-horizon forecasting requires. The subsequent sections describe the architecture of TSA-LLM, which is based on the GPT framework.
3.2.1 Input Embedding
The input embedding layer transforms the patch-structured samples ${x}^{P}_{ij}\in\mathbb{R}^{P\times L_{p}}$ into the initial hidden state $\mathbf{z}_{ij}^{0}$. This process is defined as:
$\mathbf{z}_{ij}^{0}={x}_{ij}^{P}\cdot\mathbf{W}^{p}+\mathbf{b}^{p}+\mathbf{E}_{\text{pos }}\in\mathbb{R}^{P\times d}.$ (9)
In this formulation, each patch in ${x}_{ij}^{P}$ is first mapped to an embedding of dimension $d$ via a learnable linear projection with weight matrix $\mathbf{W}^{p}\in\mathbb{R}^{L_{p}\times d}$ and bias vector $\mathbf{b}^{p}\in\mathbb{R}^{d}$. Subsequently, a positional encoding matrix, $\mathbf{E}_{\text{pos}}\in\mathbb{R}^{P\times d}$, is added to this projection, which is learned dynamically during fine-tuning to provide the model with essential information about the absolute and relative positions of each patch. This final representation, $\mathbf{z}_{ij}^{0}$, thus contains both the features of each patch and their sequential context, ready for processing by the transformer layers.
3.2.2 Transformer Block Architecture
The backbone of TSA-LLM is composed of a stack of $L$ T-blocks adapted from the GPT model, as illustrated in Figure 2(e). Each block performs a transformation that refines an input sequence $\mathbf{z}^{l-1}$ into an output sequence $\mathbf{z}^{l}$:
$\mathbf{z}^{l}=\operatorname{Transformer-Block}(\mathbf{z}^{l-1}),~\text{for }l=1,\dots,L.$ (10)
This transformation is executed via two primary sub-layers—a multi-head causal self-attention mechanism (MHCA) and a position-wise feed-forward network (FFN)—both embedded within a residual framework that utilizes skip connections and layer normalization.
The MHCA layer containing $h$ parallel causal attention heads, endows the model with the ability to dynamically focus on relevant past information, as illustrated in Figure 1(b). Within each head, $\mathbf{z}_{ij}^{0}$ is projected into queries ($\mathbf{Q}$), keys ($\mathbf{K}$), and values ($\mathbf{V}$) using (2) to compute causal attention scores:
$\text{Cal-atten}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}+\mathcal{M}_{\text{causal}}\right)\mathbf{V},$ (11)
where the lower-triangular mask, $\mathcal{M}_{\text{causal}}$ is formulated as follows:
$\mathcal{M}_{\text{causal}}(r,c)=\begin{cases}0&\text{if }r\geq c,\\ -\infty&\text{otherwise}.\end{cases}$ (12)
Pivotal to this operation is the causal mask, $\mathcal{M}_{\text{causal}}$, which nullifies the influence of future tokens by setting their corresponding attention scores to -$\infty$. This architectural constraint is what makes the model inherently suitable for iterative, generative tasks, as it forces the model to generate predictions based solely on available historical data, mimicking online dynamics prediction [liu2024timer]. The outputs of the $h$ heads are subsequently concatenated and linearly transformed by $\mathbf{W}^{O}$ to yield the multi-head representation as shown in (13).
$\text{MHCA}(\mathbf{z}_{ij}^{0})=\text{Concat}(\text{Cal-atten}_{1},\dots,\text{Cal-atten}_{h})\mathbf{W}^{O}$ (13)
The output of the multi-head attention sub-layer, after a residual connection and layer normalization, is then passed through a position-wise FNN, formulated as follows:
$\begin{aligned} &\mathbf{h}_{\mathrm{att}}=\operatorname{LayerNorm}\left(\operatorname{MHCA}\left(\textbf{z}_{ij}^{0}\right)+\textbf{z}_{ij}^{0}\right),\\ &\textbf{z}_{ij}^{1}=\operatorname{LayerNorm}\left(\operatorname{FFN}\left(\mathbf{h}_{\mathrm{att}}\right)+\mathbf{h}_{\mathrm{att}}\right),\end{aligned}$ (14)
where the FNN module introduces nonlinearity through a two-layer fully connected network:
$\operatorname{FFN}(\mathbf{h}_{\mathrm{att}})=\operatorname{GELU}(\mathbf{\mathbf{h}_{\mathrm{att}}}\cdot\mathbf{W}_{1}+\mathbf{b}_{1})\cdot\mathbf{W}_{2}+\mathbf{b}_{2}.$ (15)
In (15), $\mathbf{W}_{1}\in\mathbb{R}^{d\times d_{\mathrm{ff}}}$, $\mathbf{W}_{{2}}\in\mathbb{R}^{d_{\mathrm{ff}}\times d}$ are learnable parameters and $d_{\mathrm{ff}}$ is the hidden dimension. $\operatorname{GELU(\cdot)}$ activation provides a smooth and non-linear transformation, while the final projection $\textbf{W}_{2}$ maps features back to the input dimensionality $d$, facilitating the composition of deep transformer architectures through block stacking.
3.2.3 Output Projection
The output layer transforms the final latent representation from the transformer, $\mathbf{z}_{ij}^{L}\in\mathbb{R}^{P\times d}$, into the prediction sequence, as illustrated in Figure 2(f). The process begins by flattening the patch-level representations in a single vector ${z}_{ij}^{\text{flat}}=\operatorname{Flatten}({z}_{ij}^{L})\in\mathbb{R}^{Pd}$ to consolidate all temporal features. This flattened vector is then mapped to the desired prediction horizon, via a learnable linear projection:
${x}_{ij}^{\text{pred}}={z}_{ij}^{\text{flat}}\cdot\mathbf{W}^{\text{out}}+\mathbf{b}^{\text{out}}\in\mathbb{R}^{L_{\text{pred}}},$ (16)
where $\mathbf{W}^{\text{out}}\in\mathbb{R}^{Pd\times L_{\text{pred}}}$ and $\mathbf{b}^{\text{out}}\in\mathbb{R}^{L_{\text{pred}}}$ are the learnable weights and biases. To restore the prediction to its original physical scale, an inverse normalization operation is applied:
$\hat{x}_{ij,\text{target}}=\sigma_{ij}\cdot{x}_{ij}^{\text{pred}}+\mu_{ij},$ (17)
where the mean ($\mu_{ij}$) and standard deviation ($\sigma_{ij}$) are the same values stored from the input sample’s normalization step (see (6)). The resulting vector, $\hat{x}_{ij,\text{target}}\in\mathbb{R}^{L_{\text{pred}}}$, represents the final physical prediction for the target sequence of length.
3.3 Two-Stage Fine-tune Strategy
3.3.1 Freeze-and-Finetune Strategy
To balance adaptation to the TSA task with the preservation of knowledge from pre-trained T-blocks, we propose a parameter-efficient fine-tuning strategy. In this approach, the core T-blocks are kept frozen, while only the parameters of task-specific components, namely the input embedding layer, the layer normalization, and the output projection head, are updated during fine-tuning as illustrated in Figure 2. This strategy is designed to preserve the robust, generic feature extraction capabilities that the T-blocks acquire during pre-training. As detailed in Appendix A, the T-blocks are inherently optimized to find self- and cross-alignments within input sequences during pre-training stage, making it a powerful, modality-agnostic sequence feature extractor. Freezing the T-blocks leverages this proven capability. The efficacy of this strategy is validated in our case study. The high feature stability and co-direction (detailed in Section 4.8) confirm that the generic features extracted by the pre-trained T-blocks are highly effective for power system dynamics sequences. The following sections detail the proposed two-stage fine-tuning approach, where TeaF trains on sequence segments with ground truth sequences as input, and SchS addresses trajectory-level learning with probabilistically mixed inputs of model predictions and ground truth.
1:Input: TSA-LLM $\mathbf{F}$, Training set $D_{\text{train}}$, Epochs $E_{\text{TeaF}}$,
Learning rate $\alpha$, Integer $K$
2:Initialize: Optimizer for $\mathbf{F}$ with learning rate $\alpha$
3:for $\text{k}=1,2,\ldots,E_{\text{TeaF}}$ do
4:  for each sample $(x_{ij,\text{input}},x_{ij,\text{target}})\in D_{\text{train}}$ do
5:   $\hat{x}_{ij,\text{target}}\leftarrow\mathbf{F}(x_{ij,\text{input}})$
6:   $\mathcal{L}_{\text{TeaF}}\leftarrow\text{MSE}(\hat{x}_{ij,\text{target}},x_{ij,\text{target}})$
7:   Perform back-propagation on $\mathcal{L}_{\text{TeaF}}$ and update $\mathbf{F}$
8:  end for
9:end for
10: Initialize empty list $L_{\text{errors}}$
11:for each trajectory $x_{i}\in D_{\text{train}}$ do
12:  $\hat{x}_{i}\leftarrow\text{IterativePredict}(\mathbf{F},x_{i})$
13:  $e\leftarrow\text{MSE}(\hat{x}_{i},x_{i})$, Add $(e,x_{i})$ to $L_{\text{errors}}$
14:end for
15: Sort $L_{\text{errors}}$ by error in descending order
[…]
17:Output: Pre-trained model $\mathbf{F}$, Hard-case set $D_{\text{hard}}$
Algorithm 1 Teacher Forcing Fine-Tuning
3.3.2 Teacher Forcing Stage
During the first TeaF stage, the proposed TSA-LLM, represented by $\operatorname{\mathbf{F}}(\cdot)$, is conditioned on the ground-truth input segments, $x_{ij,\text{input}}$, to generate a prediction for the target sequence (denoted as $\hat{{x}}_{ij,\text{target}}=\operatorname{\mathbf{F}}({x}_{ij,\text{input}})$). The model parameters are optimized by minimizing the mean squared error (MSE) loss, $\mathcal{L}_{\text{TeaF}}$, between the predicted output and the ground-truth target:
$\mathcal{L}_{\text{TeaF}}=\frac{1}{L_{\text{pred}}}\sum_{t=1}^{L_{\text{pred}}}\left(\hat{x}_{ij,\text{target}}(t)-x_{ij,\text{target}}(t)\right)^{2}.$ (18)
This TeaF stage, as presented in Algorithm 1, provides a robust model initialization that accelerates convergence in subsequent SchS phases while enabling the model to achieve high fidelity on less complex transient trajectories. It is noteworthy that for numerous existing approaches, their entire training process consists solely of this TeaF paradigm [zhao2022structure, ye2024use, liu2024timer]. However, the iterative prediction process employed in online TSA application, where the TSA-LLM recursively conditions on its own outputs as illustrated in Figure 1(c), presents a training-inference mismatch. The model, trained via TeaF on ground-truth data, suffers from this mismatch and is not equipped to handle its own prediction errors, leading to cumulative error propagation in complex unstable trajectories [lamb2016professor].
3.3.3 Schedule Sampling Phase
To mitigate the previously described training-inference mismatch, a scheduled sampling fine-tuning stage is proposed. Following a curriculum learning strategy [lamb2016professor], SchS focuses on the top-$K$ error-prone trajectories, denoted by $x_{i}(t)\in D_{\text{hard}}$ for $t\in[1,T]$ identified within $D_{\text{train}}$, as identified in Algorithm 1. Critically, by emulating the iterative, self-conditioned prediction of online deployment, this process compels the model to develop robustness against its own error accumulation, thereby improving long-horizon iterative prediction in online TSA tasks.
During SchS stage, given an initial input sequence $x_{i,\text{input}}\in\mathbb{R}^{L_{\text{seq}}}$ spanning time steps $1:L_{\text{seq}}]$ from trajectory $x_{i}$, the model first generates a prediction for the next time window $L_{\text{seq}}+1,L_{\text{seq}}+L_{\text{pred}}]$:
$\hat{\mathbf{x}}_{i,\text{target}}=\mathbf{F}\left(x_{i,\text{input}}\right)\in\mathbb{R}^{L_{\text{pred}}}.$ (19)
To generate the subsequent prediction, a new input sequence, $x_{i,\text{input}}^{\text{next}}$, is constructed by sliding the window forward. The core of the SchS strategy lies in how this new input is assembled. It consists of a known segment from the original ground-truth trajectory ($x_{i,\text{orig}}$) and a ”mixed” segment ($x_{i,\text{mixed}}$), which is probabilistically sampled from either the ground-truth target ($x_{i,\text{target}}$) or the model’s own previous prediction ($\hat{x}_{i,\text{target}}$):
$x_{i,\text{mixed}}=\begin{cases}x_{i,\text{target}},&\text{with probability }\varepsilon_{k}\\ \hat{x}_{i,\text{target}},&\text{with probability }1-\varepsilon_{k}\end{cases}$ (20)
where $\varepsilon_{k}$ is the sampling rate for the current epoch $k$. The new input for the model is then a concatenation of these segments:
$x_{i,\text{input}}^{\text{next}}=[x_{i,\text{orig}},x_{i,\text{mixed}}].$ (21)
This probabilistic mixing is governed by a sampling rate $\varepsilon_{k}$ that follows a linear decay schedule across training epochs. This gradually transitions the model from teacher forcing ($\varepsilon_{k}=1$) to a fully iterative mode ($\varepsilon_{k}=0$):
$\varepsilon_{k}=\begin{cases}1,&\text{ if }k<E_{\text{start}}\\ 1-\frac{k-E_{\text{start}}}{E_{\text{max}}-E_{\text{start}}},&\text{ otherwise }\end{cases}$ (22)
where $E_{\text{start}}$ is the epoch at which decay begins, and $E_{\text{max}}$ is the maximum training epoch in SchS stage. The fine-tuning loss, $\mathcal{L}_{\mathrm{SchS}}$, aggregates the MSE across the predicted trajectory sequence $\hat{x}_{i}(t)$:
$\mathcal{L}_{\mathrm{SchS}}=\frac{1}{(T-L_{\mathrm{seq}})}\sum_{t=L_{\mathrm{seq}}+1}^{T}\left(\hat{x}_{i}(t)-x_{i}(t)\right)^{2}.$ (23)
The complete SchS procedure is detailed in Algorithm 2.
1:Input: Pre-trained model $\mathbf{F}$, Hard-case set $D_{\text{hard}}$, Epochs $E_{\text{max}}$, Scheduled epoch $E_{\text{start}}$, Learning rate $\alpha$
2:Initialize: Optimizer for $\mathbf{F}$ with learning rate $\alpha$
3:for $k=1,2,\ldots,E_{\text{max}}$ do
4:  Calculate sampling probability $\varepsilon_{k}$ using Eq. (22)
5:  for each trajectory $x_{i}\in D_{\text{hard}}$ do
6:   Initialize: Total loss $\mathcal{L}_{\text{SchS}}\leftarrow 0$
7:   $x_{i,\text{input}}\leftarrow x_{i}(1:L_{\text{seq}})$
8:   for $j=1,2,\ldots,\lfloor(T-L_{\text{seq}})/L_{\text{pred}}\rfloor$ do
9:    $\hat{x}_{i,\text{target}}\leftarrow\mathbf{F}(x_{i,\text{input}})$
10:    $\mathcal{L}_{\text{SchS}}\leftarrow\mathcal{L}_{\text{SchS}}+\text{MSE}(\hat{x}_{i,\text{target}},x_{i,\text{target}})$
11:    $x_{i,\text{input}}^{\text{next}}\leftarrow$ Construct next input with $x_{i,\text{input}}$, $\hat{x}_{i,\text{target}}$,
${x}_{i,\text{target}}$, and $\varepsilon_{k}$ based on (20) and (21)
12:    $x_{i,\text{input}}\leftarrow x_{i,\text{input}}^{\text{next}}$
13:   end for
14:   Perform back-propagation on $\mathcal{L}_{\text{SchS}}$
15:   Update model $\mathbf{F}$ parameters
16:  end for
17:end for
18:Output: Final fine-tuned TSA-LLM
Algorithm 2 Scheduled Sampling Fine-Tuning
3.4 Online Application
While channel independence processes each of the $n_{x}$ state variables separately in the offline fine-tuning stage, we mitigate potential computational overhead in online applications by treating the $n_{x}$ channels as a unified mini-batch to exploit GPU parallelization and concurrent predict of all state variables. The online prediction begins with an initial multivariate observation from PMUs, denoted as $\textbf{x}^{(0)}_{\text{input}}\in\mathbb{R}^{n_{\textbf{x}}\times L_{\text{seq}}}$. This multi-channel input is decomposed into $n_{x}$ independent univariate patch sequences $\mathbf{x}_{\text{patch}}^{(0)}\in\mathbb{R}^{P\times L_{P}}$ as described in subsection 3.1, which are then stacked along the batch dimension via $\operatorname{Batch}(\cdot)$, forming $\textbf{x}^{(0)}_{\text{batch}}\in\mathbb{R}^{n_{\textbf{x}}\times P\times L_{p}}$. The TSA-LLM then predicts the target sequence of all state variables $\hat{\textbf{x}}^{(0)}_{\text{target}}\in\mathbb{R}^{n_{\textbf{x}}\times L_{\text{pred}}\times 1}$. Then, iterative prediction for each step $j$ is performed sequentially as follows:
[…]
where $\oplus$ denotes the sliding window update mechanism, which discards the oldest input segment and appends the latest prediction, and $\operatorname{Preprocess}(\cdot)$ follows the data processing pipeline in subsection 3.1. This fully iterative online prediction aligns with the SchS strategy where the sampling probability $\epsilon_{k}$fixed at 0, forcing the model to rely entirely on its own generated trajectory.
3.5 Performance Indices Evaluation
To provide a comprehensive analysis, the model’s performance is evaluated on three distinct stability situations. The stable (S) situation comprises trajectories exclusively from stable OCs, while the unstable (U) situation contains only those from unstable OCs. A third, hybrid (H) situation is composed of a combination of both stable and unstable trajectories to assess overall generalization. The performance metrics ($MAE$, MSEMSE) for each category $\mathrm{C}\in\{\mathrm{S},\mathrm{U},\mathrm{H}\}$ are formally defined by averaging the error over all trajectories in the corresponding stability situations:
${MAE}_{\mathrm{C}}$ $=\sum_{x\in\mathcal{D}_{\mathrm{C}}}\sum_{i=1}^{n_{x}}\sum_{t=L_{\text{seq }}+1}^{T}\frac{\left|\hat{x}_{i}(t)-x_{i}(t)\right|}{|{D}_{\mathrm{C}}|\cdot n_{x}\cdot\left(T-L_{\text{seq}}\right)},$ (25)
${MSE}_{\mathrm{C}}$ $=\sum_{x\in\mathcal{D}_{\mathrm{C}}}\sum_{i=1}^{n_{x}}\sum_{t=L_{\text{seq }}+1}^{T}\frac{\left(\hat{x}_{i}(t)-x_{i}(t)\right)^{2}}{|{D}_{\mathrm{C}}|\cdot n_{x}\cdot\left(T-L_{\text{seq}}\right)},$
where ${D}_{\mathrm{C}}$ is the test set for category C, $|{D}_{\mathrm{C}}|$ is the number of trajectories within it, and $T_{\text{pred}}=T-L_{\text{seq}}$ is the prediction length. The terms $\hat{x}_{i}(t)$ and $x_{i}(t)$ denote the predicted and ground-truth values, respectively, for the $i$-th dynamic trajectories at time $t$. This categorized analysis allows us to not only quantify but also explain the performance disparities observed in its handling of stable versus unstable system dynamics.
4 Case study
The proposed TSA-LLM framework is rigorously evaluated on the New England 39-bus [zhao2022structure] and modified Iceland 189-bus [cai2020data] systems. On the 39-bus system, we assess the framework’s zero-shot generalization to mixed-stability OCs and unseen faults and conduct a representation analysis of its unstable OC handling mechanism. The 189-bus system is applied to validate the model’s few-shot scalability on a more complex, heterogeneous grid. The evaluation is completed with analyses of training/inference costs, an ablation study of model components, and a final qualitative demonstration of the predictive enhancements gained from using frozen T-blocks.
4.1 Data Generation and Model Settings
4.1.1 Test System and Simulation Settings
Two standard test systems were applied: the New England 39-bus system (10 SGs, 34 lines) and a modified Iceland 189-bus system (33 SGs, 206 lines). The latter was adapted for 20% RES penetration by replacing two SGs with wind units whose speeds were sampled from a Weibull distribution [lu2024advanced]. Simulations includes N1N-1, N2N-2 and N3N-3 faults for the 39-bus system and N1N-1 faults for the 189-bus system. For each contingency, simulations are conducted as follows: 1) All system loads were randomly varied between 70% and 130% of their nominal values; 2) An optimal power flow was solved using MATPOWER to establish the pre-fault OCs; 3) All SGs were represented by the fourth-order model; 4) Each contingency was simulated by applying the specified number of concurrent three-phase-to-ground faults to random lines at t=1.0st=1.0~s and cleared after a random duration of up to 0.3 s; 5) The post-fault dynamic response was simulated for 10 seconds with a 0.02 s time step by Power System Analysis Toolbox.
4.1.2 Dataset Organization
For fine-tuning and validation, a primary dataset of 9,000 trajectories (rotor angles and speeds) was generated from N1N-1 contingencies on the 39-bus system, covering a mix of stable and unstable trajectories. This was partitioned into 8,000 training ($D_{\text{train}}$) and 1,000 validation ($D_{\text{val}}$) trajectories. Each trajectory was then segmented via a sliding window ($L_{\text{seq}}=65,L_{\text{pred}}=1$) to yield approximately $1.8\times 10^{7}$ training and $2\times 10^{6}$ validation samples. Moreover, a 3,300-trajectory set ($D_{\text{train}}^{189\text{bus}}$) generated from the 189-bus system was used for cross-system few-shot fine-tuning. Model performance was evaluated against four distinct test sets, each designed to assess a specific aspect of universality. The details of these datasets are summarized in Table 2. The subsequent analysis primarily presents rotor angle predictions and results for other variables like rotor speeds are analogous and detailed in subsection 4.6.
Table 2: Test Sets Overview
Test Set Tested Universality Test Mode System Fault Trajectories
$D_{\text{test}}^{N-1}$ Mixed OCs Zero-shot 39-bus N1N-1 1000
$D_{\text{test}}^{N-2}$ Unseen faults Zero-shot 39-bus N2N-2 1,000
$D_{\text{test}}^{N-3}$ Unseen faults Zero-shot 39-bus N3N-3 1,000
$D_{\text{test}}^{189\text{bus}}$ Heterogeneous system Few-shot 189-bus N1N-1 3,300
4.1.3 Model Implementation Details
The proposed TSA-LLM, based on the GPT architecture (124M parameters, 12 layers, 12 heads, 768-dim embedding), was fine-tuned for up to 10 epochs using a two-stage TeaF and SchS scheme. Training utilized the Adam optimizer with a $1\times 10^{-4}$ initial learning rate managed by a cosine annealing scheduler and an early stopping criterion. To evaluate its performance, TSA-LLM was benchmarked against three baseline models. The first two are the DNR framework from [zhao2022structure] and a standard LSTM network as presented in [ye2024use], for which we employed the hyperparameter settings as reported in their respective publications. The third baseline is an Encoder-only Transformer (ENC), constructed based on [vaswani2017attention] to highlight performance differences arising from different attention mechanisms. This ENC model employs the bidirectional attention mechanism and is configured with 3 transformer layers, 6 bidirectional attention heads, and a 128-dim embedding. All baselines were implemented for multivariate dynamics prediction in PyTorch and trained on NVIDIA A100 GPU.
4.2 Evaluation on Mixed Stability OCs
4.2.1 Performance Comparison
This section evaluates the model’s robustness against mixed stable or unstable OCs on the $D_{\text{test}}^{N-1}$, which is collected in $\{\text{S},\text{U},\text{H}\}$ situations as discussed in subsection 3.5. The quantitative results are summarized in Table 3, where bold and underlined values denote the best and second-best performance, respectively. As shown, TSA-LLM substantially outperforms all baseline models across nearly all metrics, with the performance gap being most pronounced in challenging unstable and hybrid cases. While traditional methods like DNR and LSTM struggle significantly in these scenarios with DNR’s $MSE_{U}$ value even exceeding 100, TSA-LLM achieves remarkable error reductions. Specifically, TSA-LLM reduces $MAE_{U}$, $MSE_{U}$, $MAE_{H}$, and $MSE_{H}$ by at least 84.49%, 97.27%, 83.49%, and 97.68%, respectively, compared to baselines. These quantitative findings are further corroborated by visualization of the rotor angle trajectories predicted by TSA-LLM in Figure 3, which show a close tracking of the ground truth. This significant disparity in performance can be attributed to fundamental architectural differences. Traditional RNN-based models (LSTM and DNR) suffer from vanishing gradient issue and exhibit constrained model capacity, which fundamentally limits their ability to capture complex nonlinear dynamics in unstable trajectory prediction. In contrast, the transformer architecture, utilized by both our proposed TSA-LLM and the ENC baseline, is far more effective at capturing long-range temporal dependencies, resulting in demonstrably superior accuracy in these challenging scenarios.
4.2.2 Explaining Robust Prediction in Unstable Scenarios
To further investigate the source of TSA-LLM’s robustness, we conducted an interpretive analysis focused on its feature representations of OOS dynamics, which are critical in unstable scenarios. We selected a typical unstable case from the 39-bus system where SG $\delta_{5}$ loses synchronism (Figure 4(a)). Final-layer feature representations for the OOS SG ($\delta_{5}$) and two stable SGs ($\delta_{1},\delta_{10}$) were then extracted and projected into 2D space from both TSA-LLM and an identically trained channel-independent LSTM model for visualization with t-SNE [zhou2023one]. The t-SNE visualizations reveal a stark contrast between the two architectures. For TSA-LLM (Figure 4(b)), features from $\delta_{5}$ form a distinct and clearly separable cluster, indicating effective learning of instability characteristics. However, the LSTM model exhibits substantial mixing of these representations (Figure 4(c)). This demonstrates the inherent difficulty of RNN architectures in distinguishing the unique behavior of OOS generators, which correlates with their lower prediction accuracy and underscores the superior feature learning of the proposed attention-based architecture.
Table 3: New England 39-Bus System: Evaluation Results on Mixed Stability OCs
Model $MAE_{S}$ $MSE_{S}$ $MAE_{U}$ $MSE_{U}$ $MAE_{H}$ $MSE_{H}$
LSTM 0.120 0.282 2.675 19.269 0.946 5.932
DNR 0.128 0.200 7.073 133.614 1.934 35.367
ENC 0.059 0.006 0.557 0.703 0.284 0.360
TSA-LLM 0.055 0.007 0.415 0.525 0.156 0.137
(a) Actual $\delta$ in stable OC
(b) Actual $\delta$ in unstble OC
(c) Predicted $\delta$ in stable OC
(d) Predicted $\delta$ in unstble OC
Figure 3: New England 39-bus system: The evaluation of rotor angle under the stable and unstable OC. The predicted $\delta$ by TSA-LLM ((c) and (d)) demonstrates close alignment with ground truth ((a) and (b)).
Figure 4: The t-SNE visualization of stable/unstable sample feature maps for the proposed TSA-LLM and LSTM.
4.3 Zero-shot Generalizability on Unseen Faults
This section evaluates the model’s zero-shot generalizability to unseen, severe contingencies (N2N-2 and N3N-3) under mixed stability conditions. As detailed in Table 4 and Table 5, all baseline models exhibited a significant performance collapse when confronted with unseen faults. The $MSE_{H}$ for the RNN-based LSTM and DNR models surged by factors of approximately 13 and 30, respectively, compared to their N1N-1 performance. Even the transformer-based ENC model underperformed, with its $MSE_{U}$ exceeding 1.5. In contrast, TSA-LLM demonstrated superior zero-shot generalization, with its error on unstable samples ($MAE_{U},MSE_{U}$) remaining remarkably stable and increasing by less than 0.01 relative to the N1N-1 scenario. This robustness provides a substantial advantage, with TSA-LLM reducing $MSE_{H}$ by nearly 71.05%–99.17% for N2N-2 settings and 84.27%–99.78% for N3N-3 settings relative to the baselines. These performance disparities are rooted in fundamental architectural and training differences. The failure of DNR, for instance, stems not only from the inherent modeling constraints of its RNN architecture but also from its lack of a mechanism to suppress error propagation, leading to severe error accumulation in complex and unstable scenarios. TSA-LLM incorporates a TeaF and SchS scheme specifically to build resilience to its own predictive errors, ensuring stable long-horizon forecasting. Furthermore, when compared to the ENC baseline, TSA-LLM’s advantage is twofold. Its larger model scale provides greater expressive capacity, while its causal attention mechanism is better aligned with the nature of online TSA tasks than the ENC’s bidirectional attention, explaining its superior performance.
Table 4: New England 39-Bus System: Zero-shot Prediction Capacity Comparison under N-2 Setting
Model $MAE_{S}$ $MSE_{S}$ $MAE_{U}$ $MSE_{U}$ $MAE_{H}$ $MSE_{H}$
LSTM 0.152 0.736 2.663 21.880 1.379 19.708
DNR 0.200 0.760 9.193 140.934 2.647 66.703
ENC 0.083 0.010 0.793 1.593 0.781 0.563
TSA-LLM 0.042 0.003 0.466 0.498 0.168 0.163
Table 5: New England 39-Bus System: Zero-shot Prediction Capacity Comparison under N-3 Setting
Model $MAE_{S}$ $MSE_{S}$ $MAE_{U}$ $MSE_{U}$ $MAE_{H}$ $MSE_{H}$
LSTM 0.082 0.610 2.812 27.246 1.092 15.837
DNR 0.092 0.020 5.695 142.549 2.144 70.485
ENC 0.031 0.086 0.864 1.567 0.671 0.981
TSA-LLM 0.029 0.003 0.457 0.566 0.180 0.154
4.4 Scalability on heterogeneous system
This section evaluates TSA-LLM’s ability to generalize and adapt to a heterogeneous power system with limited target-system data. The model’s inherent cross-system potential was initially evaluated through a direct zero-shot test on the $D_{\text{test}}^{189\text{bus}}$ dataset, as its channel independence design permits direct application to systems of varying dimensions. Despite the significant topological and dynamic differences between the systems, TSA-LLM achieved an impressive $MSE_{H}$ of approximately 0.752. This zero-shot performance is remarkably strong, even surpassing the in-distribution performance of baseline models like LSTM and DNR on their native 39-bus system as shown in Table 3.
Furthermore, to assess the model’s few-shot learning capability, TSA-LLM pre-trained on the 39-bus system was further fine-tuned using varying proportions of $D_{\text{train}}^{189\text{bus}}$, employing the proposed TeaF and SchS schemes. Performance was subsequently evaluated on $D_{\text{test}}^{189\text{bus}}$. As shown in Figure 5, the $MSE_{H}$ of TSA-LLM rapidly deceases with increasing proportions of the 189-bus fine-tuning data and trends for other metrics are generally consistent. For a stringent comparison, the ENC model was trained on the full $D_{\text{train}}^{189\text{bus}}$ dataset to serve as an ”expert” baseline for the target system. The results demonstrate that TSA-LLM achieves performance comparable to this fully-trained expert after fine-tuning on only 4%–5% of the available target-system data. This remarkable data efficiency is visually confirmed by the predicted trajectories from the TSA-LLM fine-tuned by 5% samples from $D_{\text{train}}^{189\text{bus}}$ in Figure 5. Furthermore, when fine-tuned on the complete 189-bus dataset, TSA-LLM reduces the $MSE_{H}$ by nearly 89.62% compared to the expert ENC model, underscoring its superior architectural capacity. This cross-system performance introduces the potential to replace the costly, expert-intensive process of building custom-built models for large-scale systems by enabling an efficient methodology where a foundational model is trained on smaller systems and subsequently adapted with minimal target-system data. Beyond this immediate economic advantage, our findings reveal a clear path for true performance scalability. The model is designed to continually enhance its multi-system, multi-scenario proficiency by incorporating data from an ever-expanding corpus of power systems, paving the way for a single, continually evolving foundation TSA model.
(a) Few-shot scalability
(b) Actual $\delta$ in stable OC
(c) Actual $\delta$ in unstable OC
(d) Predicted $\delta$ in stable OC
(e) Predicted $\delta$ in unstable OC
Figure 5: Iceland 189-bus system: few shot scalability. To facilitate presentation, prediction results for a subset of 15 SGs are displayed.
4.5 Ablation Study
To assess the individual contributions of the dual-phase training strategy (TeaF and SchS) and the temporal patch, a series of ablation studies were performed, with results summarized in Table 6. Model variants, denoted ”w/o” for the exclusion of a component, were trained on the $D_{\text{train}}$ and $D_{\text{val}}$. Their performance was subsequently evaluated across all test sets ($D_{\text{test}}^{N-1}$, $D_{\text{test}}^{N-2}$, $D_{\text{test}}^{N-3}$, and $D_{\text{test}}^{189\text{bus}}$), with a 5% fine-tuning step on $D_{\text{train}}^{189\text{bus}}$ applied specifically for the final cross-system test. The results confirm that all components are integral to the model’s performance. The removal of TeaF and SchS, which form the dual-phase training strategy, led to the most significant performance degradation, with average error increases of nearly 68.76% and 41.16%, respectively, highlighting their importance in robustly modeling dynamics and suppressing cumulative errors. Lastly, removing the patch operation resulted in a nearly 21.26% performance decline, underscoring its importance in helping the attention mechanism model rich and informative local dynamic patterns.
Table 6: Ablation Study Comparing the Performance of Full Proposed Model against Variants with Removed Components.
Model $D_{\text{test}}^{N-1}$ $D_{\text{test}}^{N-2}$
$MAE_{H}$ $MSE_{H}$ $MAE_{H}$ $MSE_{H}$
w/o Teacher Forcing 0.194 0.296 0.241 0.329
w/o Schedule Sampling 0.189 0.178 0.192 0.258
w/o Patch 0.161 0.145 0.192 0.230
TSA-LLM 0.156 0.137 0.168 0.163
Model $D_{\text{test}}^{N-3}$ $D_{\text{test}}^{189bus}(5\%)$
$MAE_{H}$ $MSE_{H}$ $MAE_{H}$ $MSE_{H}$
w/o Teacher Forcing 0.310 0.222 0.113 0.299
w/o Schedule Sampling 0.267 0.202 0.138 0.210
w/o Patch 0.198 0.173 0.114 0.196
TSA-LLM 0.180 0.154 0.072 0.156
4.6 Validation on Angular Speed
Recognizing the operational need for comprehensive state variable monitoring, this subsection evaluates the proposed model’s predictive performance for rotor angular speed, with all predictions presented in per-unit (p.u.) values. As summarized in Table 7, the $MSE_{H}$ results indicate that the proposed model demonstrates an even more substantial performance advantage in rotor speed prediction compared to its rotor angle prediction capabilities. Across the four primary evaluation tasks, the proposed model achieves performance improvements of up to two orders of magnitude over baselines. For instance, when trained with the complete $D_{\text{train}}^{\text{189bus}}$, the proposed approach reduced the $MSE_{H}$ by nearly 95.95% relative to the leading baseline model (ENC) on $D_{\text{test}}^{\text{189bus}}$. Furthermore, in a few-shot learning scenario on this system, the proposed method matched the performance of the fully trained ENC model while utilizing only approximately 3% of the training data—a smaller proportion than required for the rotor angle prediction task. These findings underscore the proposed method’s efficacy and versatility in predicting various critical state variables.
Table 7: Angular Speed Prediction Performance Comparison by $MSE_{H}$
Model $D_{\text{test}}^{N-1}$ $D_{\text{test}}^{N-2}$ $D_{\text{test}}^{N-3}$ $D_{\text{test}}^{\text{189bus}}(100\%)$
LSTM 0.0827 0.0904 0.1002 0.2036
DNR 0.0395 0.0457 0.0389 0.1235
ENC 0.0427 0.0478 0.0529 0.1136
TSA-LLM 0.0009 0.0007 0.0019 0.0046 (3%)
4.7 Training and Inference Cost Comparison
This section analyzes the computational efficiency of the proposed TSA-LLM, comparing it against the ENC baseline and the traditional TDS on the large-scale 189-bus system. We assess efficiency using two key metrics: 1) the training cost per step, defined as a single forward and backward propagation pass on one batch, and 2) the inference time to predict the complete trajectories for a single operating condition (OC). The detailed results of this comparison are presented in Table 8.
4.7.1 Training Efficiency
The remarkable training efficiency of the large-scale TSA-LLM, which remains comparable to much smaller models, is almost entirely attributable to its parameter-efficient fine-tuning approach. The strategy of freezing the main T-blocks and training only 0.69% of the total parameters significantly reduces the computational complexity associated with the most intensive phase of training, namely the backward pass and parameter update. This allows the model to be fine-tuned with exceptional speed, evidenced by a per-step training time on the large 189-bus system that is only negligibly slower (0.004s) than the compact ENC baseline.
4.7.2 Inference Speed
In the inference stage, TSA-LLM exhibits a significant speed advantage, proving faster than even the smaller ENC baseline on the 189-bus system. This high efficiency stems from three core design principles. First, the patch and stride operations strategically reduce the input sequence length for the attention mechanism, which is instrumental in mitigating the quadratic computational complexity ($\mathcal{O}(N^{2})$) inherent in long-sequence modeling. Second, the causal attention mechanism results in a sparse attention weight matrix where most elements are zero, thereby simplifying the computational process of aggregating patch representations. Most critically, the online application design reframes the problem by transforming all $n_{x}$ system channels into the batch dimension. This enables the model to process the entire dynamic trajectory for a given operating condition simultaneously while fully leveraging the massive parallel processing capabilities of GPUs on large batches. Consequently, TSA-LLM not only outperforms the ENC model but also achieves an inference speedup of approximately 96.88% over traditional TDS. These results underscore the framework’s strong potential for efficient, real-time trajectory prediction.
Table 8: Training Parameters and Training/Inference Cost Comparison in the Modified Iceland 189-Bus System
Method Trainable Params Trainable Params Percentages Training (s/step) Inference (s/OC)
TDS - - - 18.25
ENC 9.6M 100% 0.034 0.59
TSA-LLM 858K 0.69% 0.038 0.57
4.8 Generic Validation of Pre-trained Parameters
4.8.1 Validation of Performance Gain from Pre-training
The remarkable universality of TSA-LLM is critically dependent on the knowledge embedded in its pre-trained T-blocks. We verified this through a direct comparison on the $D_{\text{test}}^{N-1}$ dataset between TSA-LLM and a variant whose parameters within T-blocks were randomly sampled from a standard distribution. As presented in Table 9, the model without pre-trained parameters performs poorly, whereas TSA-LLM shows dramatic improvements, including reductions of nearly 86.23% in $MAE_{H}$ and 99.41% in $MSE_{H}$. This contrast underscores that the foundational sequence modeling capabilities of pre-trained T-blocks in GPT, though trained on text, are modality-agnostic and successfully generalize to capture the complex patterns of power system dynamics.
4.8.2 Explaining Performance Gains
As detailed in Appendix A, pre-training process optimizes T-blocks in GPT model to produce stable and co-directional feature representations. To further validate the transfer of these properties to transient dynamics, two analyses were conducted. First, feature stability was quantified by the average cosine similarity between representations of same patches from the final two transformer layers. The pre-trained model exhibited an average cross-layer similarity of 0.893, significantly higher than the 0.673 from the model without pre-trained parameters, indicating superior representational consistency through deeper layers. Second, feature co-direction was evaluated by the frequency of high cosine similarity (exceeding 0.8) representation pairs within the final layer. As shown in Table 9, the pre-trained model yielded 71.44% high-similarity pairs, substantially exceeding the 46.78% from the model without pre-trained parameters. These combined findings confirm that pre-trained T-blocks retain their generic sequence feature extraction when applied to transient dynamics, empirically explaining the efficacy of the pre-trained T-blocks in TSA tasks.
Table 9: Comparison of Model Performance with and without Pre-trained Parameters
Model $MAE_{H}$ $MSE_{H}$
w/o pre-trained param 1.134 23.220
TSA-LLM 0.156 0.137
Model Feature Feature
stability co-direction (>>0.8)
w/o pre-trained param 0.673 46.78%
TSA-LLM 0.893 71.44%
5 Conclusion
This paper introduces TSA-LLM, a universal framework that leverages a pre-trained LLM to achieve robust generalization across diverse OCs, unseen faults, and heterogeneous systems with a single architecture. This framework comprises a novel data processing pipeline featuring channel-independence, sample-wise normalization, and temporal patching to handle mixed-stability and long sequences from heterogeneous systems. Moreover, a parameter-efficient, freeze-and-finetune strategy for the augmented GPT architecture is designed to preserve generic sequence feature extraction from pre-trained T-blocks while allowing for TSA task adaptation. Furthermore, a two-stage fine-tuning scheme that combines TeaF with SchS is devised to mitigate cumulative prediction errors and ensuring long-horizon reliability. Case studies on two benchmark systems demonstrate that the proposed approach exhibits exceptional universality, low training and inference costs, and a considerable degree of interpretability, establishing it as a promising foundation for real-time TSA.
Appendix A Generic Feature Extraction of T-Blocks
The decision to freeze the T-blocks in the GPT is justified by their optimization during pre-training, which molds them into powerful, generic feature extractors. Following the theoretical framework established in [zhou2023one, kim2021lipschitz], pre-training process is driven by the minimization of the attention gradient’s upper bound ($|G|_{2}$), as formulated in (26).
$\begin{aligned} |G|_{2}&\leq|A|_{2}\sum_{i}\left(a_{i,i}+\frac{1}{2}\right)\underbrace{\left|x_{i}-\sum_{j}a_{i,j}x_{j}\right|^{2}}_{\text{Self-alignment}}+\\ &|A|_{2}\sum_{i\neq j}a_{i,j}\underbrace{\left|x_{j}-\sum_{k}a_{i,k}x_{k}\right|^{2}}_{\text{Cross-alignment}}+\frac{|A|_{2}}{2}\sum_{i}\left|x_{i}\right|^{2}.\end{aligned}$ (26)
In (26), $A$ is the attention matrix, $x_{i}$, $x_{j}$ and $x_{k}$ are the token representation and $a_{i,j}$ signifies softmax probabilities from $A$. Minimizing this bound fosters two key alignment dynamics that function as an emergent form of dimensionality reduction. The Self-alignment term promotes feature stability by encouraging each feature vector $x_{i}$ to become more similar to its own attention-weighted context ($\sum a_{i,j}x_{j}$), which stabilizes the feature as they propagate through the network. Concurrently, the Cross-alignment term promotes feature co-direction by forcing feature vectors ($x_{i},x_{j}$) with high mutual attention to align with each other, mapping correlated inputs to representation vectors with similar directions. Together, these processes compel the model to represent complex sequences using a compact and consistent set of principal features. This inherent capability for generic and modality-agnostic feature extraction, forged during pre-training, is precisely why the T-block parameters are frozen for TSA tasks.